You are given the following matrix

 "A=\left(\begin{array}{cccc}a & b & c\\ d & e & f \\g & h & i\end{array} \right)."

Which of the following is the determinant of A?


  1. "aei+bfg\u2212cdh\u2212ceg\u2212afh\u2212bdi"
  2. "aei+bfg+cdh\u2212ceg\u2212afh+bdi"
  3. "aei\u2212bfg+cdh\u2212ceg+afh\u2212bdi"
  4. "aei+bfg+cdh\u2212ceg\u2212afh\u2212bdi"
  5. "\u2212aei\u2212bfg\u2212cdh+ceg+afh+bdi"
1
Expert's answer
2021-09-20T04:51:19-0400

There is a square matrix "A=""\begin{pmatrix}
 a_{11} & a_{12} & a_{13} \\
 a_{21} & a_{22} & a_{23} \\
 a_{31} & a_{32} & a_{33}
\end{pmatrix}". By defenition "det(A)" is a sum all of products of 3 elements. The elements should be taken one at a time from each row and column. A sign of product depending on the permutation.

Take a products, in each of them indexes by rows have no inversion:

1) "a_{11}a_{22}a_{33}"

Indexes by row: 1,2,3: no inversion -- 0

Indexes by column: 1,2,3: no inversions -- 0.

0 + 0 = 0, it's even number, sign is plus.

2) "a_{12}a_{23}a_{31}"

Indexes by row: 1,2,3: no inversions -- 0

Indexes by column: 2,3,1

2 and 1 are inversed, 3 and 1 are inversed -- 2 inversions.

0 + 2 = 2, it's even number, sign is plus.

3) "a_{13}a_{21}a_{32}"

Indexes by row: 1,2,3: no inversions -- 0

Indexes by column: 3,1,2

3 and 1 are inversed, 3 and 2 are inversed -- 2 inversions.

0 + 2 = 2, it's even number, sign is plus.

4) "a_{13}a_{22}a_{31}"

Indexes by row: 1,2,3: no inversions -- 0

Indexes by column: 3,2,1

3 and 2 are inversed, 3 and 1 are inversed, 2 and 1 are inversed -- 3 inversions.

0 + 3 = 3, it's odd number, sign is minus.

5) "a_{11}a_{23}a_{32}"

Indexes by row: 1,2,3: no inversions -- 0

Indexes by column: 1,3,2

3 and 2 are inversed -- 1 inversion.

0 + 1 = 1, it's odd number, sign is minus.

6) "a_{12}a_{21}a_{33}"

Indexes by row: 1,2,3: no inversions -- 0

Indexes by column: 2,1,3

2 and 1 are inversed -- 1 inversion.

0 + 1 = 1, it's odd number, sign is minus.


So, the "det(A)=a_{11}a_{22}a_{33} + a_{12}a_{23}a_{31}+a_{13}a_{21}a_{32}-a_{13}a_{22}a_{31}-a_{11}a_{23}a_{32} \n-a_{12}a_{21}a_{33}"


If there is "A = \begin{pmatrix}
 a & b & c \\
 d & e & f \\
 g & h & i
\end{pmatrix}" , using the defenition of determinant for matrix 3x3 and calculete it:

"det(A)=aei+bfg+cdh-ceg-afh-bdi"

It is answer number 4.


Answer: 4. "aei+bfg+cdh-ceg-afh-bdi"